# Effective Whitney Stratification of Real Algebraic Varieties

### v3

@@ 4.1. Real Varieties
Let $W\subset Z$ be a pair of real algebraic varieties in $\mathbb{R}^{n}$. If the difference $Z(\mathbb{C})-W(\mathbb{C})$ is either empty or a smooth $i$-dimensional complex manifold, then $M:=(Z-W)$ is either empty or a smooth $i$-dimensional real manifold.
@@
There are two cases to consider — either the image $\iota(Z)$ lies entirely within the singular locus $Z(\mathbb{C})_{\rm sing}$, or there exists some $p\in M$ with $\iota(p)\in Z(\mathbb{C})_{\rm reg}$. In the first case, since $Z(\mathbb{C})-W(\mathbb{C})$ is smooth, we know that $Z(\mathbb{C})_{\rm sing}$ lies entirely within $W(\mathbb{C})$ and hence that $\iota(Z)\subset\iota(W)$; but since we have assumed $W\subset Z$, we must have $W=Z$, whence $M$ is empty. On the other hand, let $p$ be a point in $(Z-W)$ for which $\iota(p)$ is a smooth point of $Z(\mathbb{C})$. We may safely assume that the generating ideal of $Z$ is prime in $\mathbb{C}[x_{1},\ldots,x_{n}]$ by passing to the irreducible component which contains $\iota(p)$. It now follows from [38, Theorem 12.6.1] or [23, Theorem 2.3] that $(Z-W)$ has dimension $i$. Finally, Lemma 4.1 ensures that $(Z-W)$ is a smooth real $i$-manifold. ∎
@@
Consider a connected component $M\subset(X_{i}-X_{i-1})$, and let $V\subset X_{i}$ be the irreducible component which contains $M$. Similarly, let $W\subset X_{i}(\mathbb{C})$ be the irreducible component which contains $\iota(M)$ and define $M_{\mathbb{C}}:=W-X_{i-1}(\mathbb{C})$. We note that $\iota(M)$ forms an open subset of $M_{\mathbb{C}}$, which must in turn be an $i$-stratum of $X(\mathbb{C})$. Similarly, consider a nonempty connected $N\subset S_{j}$ with $i>j$ and analogously define $N_{\mathbb{C}}\subset X_{j}(\mathbb{C})$. We will show that the pair $(M,N)$ satisfies Condition (B). To this end, consider a point $q\in N$ along with sequences ${\left\{{p_{k}}\right\}}\subset M$ and ${\left\{{q_{k}}\right\}}\subset N$ which converge to $q$. Letting $\ell_{k}$ denote the secant line $[p_{k},q_{k}]$ and $T_{k}$ the tangent plane $T_{p_{k}}M$, we assume further that the limits $\ell=\lim\ell_{k}$ and $T=\lim T_{k}$ both exist.
Let $\ell_{k}(\mathbb{C})$ be the secant line $[\iota(p_{k}),\iota(q_{k})]$ in $\mathbb{C}^{n}$ and let $T_{k}(\mathbb{C})$ be the tangent space $T_{\iota(p_{k})}M_{\mathbb{C}}$. Since $\iota(p_{k})$ and $\iota(q_{k})$ are real points for all $k$, the the linear equations defining both $\ell_{k}(\mathbb{C})$ and $T_{k}(\mathbb{C})$ as varieties are exactly the same as those defining $\ell_{k}$ and $T_{k}$, respectively. Thus, the limits $\ell(\mathbb{C})$ and $T(\mathbb{C})$ both exist because the corresponding real limits exist – one may view these as limits of real sequences inside a complex Grassmannian – and they are defined by the same algebraic equations as their counterparts $\ell$ and $T$. By definition of secant lines, the image $\iota(w)$ of any $w\in\ell$ is a real point of $\ell(\mathbb{C})$. Since the pair $(M_{\mathbb{C}},N_{\mathbb{C}})$ satisfies Condition (B) by assumption, we know that $\ell(\mathbb{C})\subset T(\mathbb{C})$, whence $\iota(w)$ must be a real point of $T(\mathbb{C})$. Since $\iota(M)$ is an open subset of $M_{\mathbb{C}}$, we have $T_{\iota(p_{k})}M_{\mathbb{C}}=T_{\iota(p_{k})}\iota(M)$ for all $k$; and by the proof of Lemma 4.1, the real points of $T_{\iota(p_{k})}\iota(M)$ are identified with $T_{q_{k}}N$. Thus, $\iota(T)$ contains all the real points of $T(\mathbb{C})$, including $\iota(w)$. Since $\iota$ is injective, we have $w\in T$ as desired. ∎
@@
The WhitStrat algorithm performs three types of operations: ideal addition, Gröbner basis computation, and primary decomposition. Each of these operations leaves the coefficient field of all intermediate polynomials unchanged. Similarly, the WhitneyPolar algorithm only uses ideal addition and Gröbner basis computation, so the same argument applies. ∎
@@ 4.3. Dominant Morphisms between Equidimensional Varieties
It is a classical result, see e.g. [43, Chapter II, Corollary 6.7], that if $f:X\to Y$ is finite in the above sense, then it is also a proper map for any field (see Remark 2.4 for the significance of this result in our context). Over $\mathbb{K}=\mathbb{C}$, a dominant morphism is finite if and only if it is proper.
@@
then the restriction of $f$ forms a proper map $(X(\mathbb{C})-W(\mathbb{C}))\to(Y(\mathbb{C})-V(\mathbb{C}))$ — see [34, Proposition 6.1] for details. Note that the polynomials defining $V(\mathbb{C})$ and $W(\mathbb{C})$ are real; take $V$ to be the real zero set of the polynomials defining $V(\mathbb{C})$, and similarly let $W$ be the real zero set of the polynomials defining $W(\mathbb{C})$. It follows immediately that the restriction of $f$ to the difference $(X-W)$ constitutes a proper map to the difference $(Y-V)$.
@@
Since the stratification $Y_{\bullet}$ of $Y$ is subordinate to $V_{\bullet}$, for each stratum $R\subset Y$ we have that $R\subset V_{i}-V_{i-1}$ for some $i$, and also since $X_{\bullet}$ is subordinate to $W_{\bullet}$ we also have $f^{-1}(R)\subset W_{i}-W_{i-1}$. If $\dim(V_{i})=\dim(W_{i})$ holds, then the map $f:(W_{i}-W_{i-1})\to(V_{i}-V_{i-1})$ is proper. An appeal to the semialgebraic version of Thom’s first isotopy lemma [10, Theorem 1] achieves the desired result. ∎
@@ 5. Full Semialgebraic Stratifications
Here the $g_{i}$’s are a finite collection of polynomials in $\mathbb{R}[x_{1},\ldots,x_{n}]$. By convention, when the number of inequalities $k$ equals zero, we have $C=\mathbb{R}^{n}$. Thus, every algebraic variety is automatically a semialgebraic set in the above sense. The sets $X$ and $C$ are not uniquely determined for a given $B$ in general — we may, for instance, safely remove any polynomial generator $f:\mathbb{R}^{n}\to\mathbb{R}$ from the defining ideal of $X$ while adding $f\geq 0$ and $-f\geq 0$ to the inequality locus. It is therefore customary to omit $X$ entirely and simply define $B$ as the set of points which satisfy a collection of polynomial inequalities. We find it convenient to write $B=X\cap C$ here because this allows us to isolate a relevant sub-class of semialgebraic sets.
@@
The following example highlights an important class of full semialgebraic sets which arise in the Morse theoretic study of complex analytic varieties — see [20, II.2].
@@
Whitney showed in [52] that every complex analytic variety $W\subset\mathbb{C}^{n}$ admits a Whitney stratification $W_{\bullet}$ where each intermediate $W_{i}$ is a complex analytic subvariety. Associated to each $i$-stratum $M\subset W_{i}-W_{i-1}$ is a stratified space $\mathscr{L}=\mathscr{L}_{M}$, which is called the complex link of $M$ and defined as follows. Consider a point $p\in M$ and let $A\subset\mathbb{C}^{n}$ be a generic $(n-i)$-dimensional affine subspace which intersects $W$ transversely at a point $p\in M$. Let $B_{\epsilon}(p)$ denote the disk containing all points within a sufficiently small radius $\epsilon>0$ around $p$ with respect to the standard euclidean distance in $\mathbb{C}^{n}$. Finally, let $\pi:A\to\mathbb{C}$ be a generic linear form sending $p$ to zero. For some positive $\delta\ll\epsilon$, one defines
@@
Since $B$ is full, we know that the interior $C^{\circ}$ of its inequality locus is a smooth open $n$-dimensional submanifold of $\mathbb{R}^{n}$. This means any open neighborhood of $X$, which is contained in $C^{\circ}$, inside $\mathbb{R}^{n}$ is identical to the same neighborhood of $X$ inside $C^{\circ}$ and therefore, the intersections $X_{i}\cap C^{\circ}$ form a Whitney stratification $X^{\prime}_{\bullet}$ of $X\cap C^{\circ}$. Alternatively the fact that $X_{i}\cap C^{\circ}$ forms a Whitney stratification $X^{\prime}_{\bullet}$ of $X\cap C^{\circ}$ follows since $C^{\circ}$ is full dimensional, meaning the intersection $X\cap C^{\circ}$ is transverse and hence the conclusion follows because the Whitney stratification of two spaces which intersect transversely is given by their intersection, [9], see also [20, Chapter 1, §1.2, (5)].
@@
Using Theorem 5.4 in practice on a full semialgebraic set $B=X\cap C$ requires constructing a stratification of the real variety $X\cap Y_{C}$; this is somewhat unsatisfactory from an efficiency perspective, since $X$ and $Y_{C}$ may intersect far away from $B$ — returning to Example 5.3, note that both $P_{1}$ and $P_{2}$ will intersect $L_{2}$ at points far from $B$. We are not aware of any existing technique which can easily overcome such suboptimality.
6. Complexity and Performance Analysis
There is a thorough complexity analysis of WhitStrat in [26, Sec 7.2]; the complexity analysis of WTBStrat is similar. Here we derive new complexity bounds and examine the practical performance of the new WhitneyPolar algorithm presented in Section 3.
6.1. Complexity Analysis of WhitneyPolar
Here we show that the complexity of WhitneyPolar from Section 3 is singly exponential in the number of variables. The main algorithmic task is the computation of polar varieties [4, 3, 47], which reduces immediately to ideal elimination. Therefore, we will express the complexity of our algorithms in terms of the complexity of elimination.
Definition 6.1.
Given positive integers $m,n,r$ and $\delta$, let $\mathcal{E}(n,m,r,\delta)$ denote the complexity of computing the elimination ideal $I\cap\mathbb{Q}[x_{1},\ldots,x_{n}]$, where $I=\langle f_{1},f_{2},\ldots,f_{r}\rangle$ is a polynomial ideal in $\mathbb{Q}[x_{1},\ldots,x_{n},y_{1},\ldots,y_{m}]$ with $\delta:=\max_{i}{\left\{{\deg f_{i}}\right\}}$.
A variety of classical sources (such as [49, Theorem 1]) reveal that
$\mathcal{E}(n,m,r,\delta)\cong r^{\mathcal{O}(1)}\delta^{\mathcal{O}(m^{2}n)}.$
Most such bounds are based on resultant computation, but in practice it is often more effective to compute elimination ideals using Gröbner basis techniques, see [11, Chapter 3, Theorem 2]. It is shown in [5, Theorem 2] that after a general linear change of coordinates, a Gröbner basis of $I$ may be computed in a number of operations bounded by $\mathcal{O}(9r\delta^{3(n+m)})$ when $r\leq n+m$. Using either resultants or Gröbner bases, we may crudely bound $\mathcal{E}(n,m,r,\delta)\cong r^{\mathcal{O}(1)}\delta^{\mathcal{O}(n^{3}+m^{3})}$.
Remark 6.2.
In the statements of the results below, we take our varieties in an $(n-1)$ dimensional ambient space. When computing polar varieties, we perform elimination after adding one extra variable, hence obtaining a total of $n$ variables in the setup of Remark 3.4(1), or of $12(n-1)+1=2n-$ 1 variables in the setup of Remark 3.4(3). To avoid a proliferation of 1’s, we will frequently bound $n-1$ by $n$ below. Further, to be agnostic with regards to (1) vs (3) in Remark 3.4, we use a bound of $2n$ in some places.
The following result establishes that the cost of checking where Condition (B) fails via Lemma 3.5 and Theorem 3.3 is singly exponential in $n$.
Lemma 6.3.
Suppose that $X$ is a variety (in $\mathbb{R}^{n-1}$ or $\mathbb{C}^{n-1}$) whose defining ideal $I$ is generated by $r$ polynomials of degree at most $\delta$. Let $Y$ be a subvariety of $X_{\rm Sing}$ with degree bounded by $\delta^{n^{2}}$. Then the number of arithmetic operations required to find all points in $Y$ where Whitney’s Condition (B) fails with respect to $X$ (via Lemma 3.5 combined and Theorem 3.3), is bounded in $(2r+1)^{\mathcal{O}(1)}\delta^{\mathcal{O}(2n^{5})}$.
Proof.
Although we use irreducible decomposition in our presentation of WhitneyPolar for simplicity of presentation, in fact one only requires equidimensional decomposition in Lines 3, 7 and 10. The equidimensional decomposition of an ideal generated by $s$ polynomials in $k$ variables of degree at most $D$ may be computed in $\mathcal{O}(sD^{k^{2}})$, see [37, Theorem 1] or [14]. Hence, for $I$, we may bound this complexity in $\mathcal{E}(n,m,r,\delta)$.
From (7) and Remark 3.4, and recalling we can always bound the degree of a variety by that of any complete intersection containing it, Bézout’s theorem guarantees that the degree of any polar variety of $X$ is no larger than $\delta^{n^{2}}$. From Remark 3.4, we see that the number of equations defining the polar variety prior to elimination is $12r+$ 1. Since the number of operations required to compute the elimination is
$(2r+1)^{\mathcal{O}(1)}(\delta)^{n^{2}\mathcal{O}(2n^{3})}\cong(2r+1)^{\mathcal{O}(1)}(\delta)^{\mathcal{O}(2n^{5})},$
it follows that the resulting number of generators is also bounded in this order. ∎
With the result for a single pair in hand we may now state and prove our complexity result for the entire WhitneyPolar algorithm.
Theorem 6.4.
Suppose $X=\mathbf{V}_{\mathbb{C}}(f_{1},\dots,f_{r})$ is an algebraic variety (in $\mathbb{R}^{n-1}$ or $\mathbb{C}^{n-1}$) where $\deg(f_{i})\leq\delta$ and with $\mu=\dim(X_{\rm sing})$. The number of arithmetic operations required by WhitneyPolar is of the order $(2r+1)^{\mathcal{O}(1)}{\delta^{\mathcal{O}(n)}}^{2\mu+3}$.
Proof.
The degree of $X_{\rm sing}$ is bounded by $\delta^{n^{2}}$, whence the complexity of the first step of WhitneyPolar algorithm is bounded in $(2r+1)^{\mathcal{O}(1)}(\delta)^{\mathcal{O}(2n^{5})}$ by Lemma 6.3. The output stratification can have at most $\mu-1=\dim(X_{\rm sing})-1$ many levels below the singular locus; at each level, the next level is computed by calculating polar varieties on the output of previous steps. Hence, having obtained a polar variety, on the subsequent level we will perform elimination on an ideal with at most $(2r+1)^{\mathcal{O}(1)}\delta^{\mathcal{O}(2n^{5})}$ generators of degree at most $(\delta^{n^{2}})^{n^{2}}=\delta^{n^{4}}$. This gives a cost of
$(2r+1)^{\mathcal{O}(1)}\delta^{\mathcal{O}(2n^{5})}(\delta^{n^{4}})^{\mathcal{O}(2n^{3})}\cong(2r+1)^{\mathcal{O}(1)}\delta^{\mathcal{O}(n^{7})}.$
Performing equidimensional decomposition on such a polar variety would be bounded in $(2r+1)^{\mathcal{O}(1)}\delta^{\mathcal{O}(2n^{5})}(\delta^{n^{4}})^{\mathcal{O}(n^{2})}$ , and hence bounded in $(2r+1)^{\mathcal{O}(1)}(\delta)^{\mathcal{O}(n^{7})}$ as well. Starting from the singular locus and iterating this procedure at most $\mu-1$ times gives the total cost of
$(2r+1)^{\mathcal{O}(1)}\delta^{\mathcal{O}(n^{2(\mu-1)+5})}=(2r+1)^{\mathcal{O}(1)}\delta^{\mathcal{O}(n^{2\mu+3})},$
as desired. ∎
In the setting of the Theorem, suppose $r^{\mathcal{O}(1)}\leq\mathcal{O}(\delta^{n})$ (which is most often the case). The WhitStrat algorithm applied to $X$ has complexity bounded by ${\delta^{\mathcal{O}({n2^{n}})}}^{\mu}$, see [26, Theorem 7.3]. On the other hand, WhitneyPolar has complexity roughly ${\delta^{\mathcal{O}(n)}}^{2\mu+3}$, so the base of the $\mu$ exponent is singly rather than doubly exponential in the number of variables $n$. This difference is primarily due to the fact that in WhitneyPolar we only need to iterate over irreducible (or equidimensional) components, both of which have number bounded in $\mathcal{O}(\delta^{n})$. On the other hand, WhitStrat requires iterating over primary components of an ideal; these can only be bounded in $\mathcal{O}(\delta^{n2^{n}})$, see Subsection 2.1.
Remark 6.5.
The complexity estimates derived above form a (very slight) improvement over the $\delta^{\mathcal{O}(n)^{6\mu}}$ bounds of Rannou et al from [42, 45]. The key difference is that the complexity estimates of Rannou rely on asymptotically fast critical points based quantifier elimination methods of Grigoriev-Vorobjov [21] and Renegar [46], whereas our bounds use only standard elimination theory over algebraically closed fields. As noted in [26, §7.3], the asymptotically fast critical points based quantifier elimination methods of Grigoriev-Vorobjov [21] and Renegar [46] have never been implemented, despite having been extensively studied for the last 30 years, and further there is little hope they ever will be implemented. In practice, all implementations of the quantifier elimination, for example in Maple or Mathematica, rely on cylindrical algebraic decomposition which, while very powerful, is also well known as very computationally expensive both in theory and in practice.
6.2. Comparisons, Practical Considerations and Tests

### v4

@@ 4.1. Real Varieties
Let $W\subset Z$ be a pair of real algebraic varieties in $\mathbb{R}^{n}$. If the difference $Z(\mathbb{C})-W(\mathbb{C})$ is either empty or a smooth $i$-dimensional complex manifold, then M:=(ZW)M:=(Z-W) is either empty or a smooth $i$-dimensional real manifold.
@@
There are two cases to consider — either the image $\iota(Z)$ lies entirely within the singular locus $Z(\mathbb{C})_{\rm sing}$, or there exists some $p\in M$ with $\iota(p)\in Z(\mathbb{C})_{\rm reg}$. In the first case, since $Z(\mathbb{C})-W(\mathbb{C})$ is smooth, we know that $Z(\mathbb{C})_{\rm sing}$ lies entirely within $W(\mathbb{C})$ and hence that $\iota(Z)\subset\iota(W)$; but since we have assumed $W\subset Z$, we must have $W=Z$, whence $M$ is empty. On the other hand, let $p$ be a point in (ZW)(Z-W) for which $\iota(p)$ is a smooth point of $Z(\mathbb{C})$. We may safely assume that the generating ideal of $Z$ is prime in $\mathbb{C}[x_{1},\ldots,x_{n}]$ by passing to the irreducible component which contains $\iota(p)$. It now follows from [35, Theorem 12.6.1] or [20, Theorem 2.3] that (ZW)(Z-W) has dimension $i$. Finally, Lemma 4.1 ensures that (ZW)(Z-W) is a smooth real $i$-manifold. ∎
@@
Consider a connected component $M\subset(X_{i}-X_{i-1})$, and let $V\subset X_{i}$ be the irreducible component which contains $M$. Similarly, let $W\subset X_{i}(\mathbb{C})$ be the irreducible component which contains $\iota(M)$ and define $M_{\mathbb{C}}:=W-X_{i-1}(\mathbb{C})$. We note that $\iota(M)$ forms an open subset of $M_{\mathbb{C}}$, which must in turn be an $i$-stratum of $X(\mathbb{C})$. Similarly, consider a nonempty connected $N\subset(X_{j}-X_{j-1})$ with $i>j$ and analogously define $N_{\mathbb{C}}\subset X_{j}(\mathbb{C})$. We will show that the pair $(M,N)$ satisfies Condition (B). To this end, consider a point $q\in N$ along with sequences ${\left\{{p_{k}}\right\}}\subset M$ and ${\left\{{q_{k}}\right\}}\subset N$ which converge to $q$. Letting $\ell_{k}$ denote the secant line $p_{k},q_{k}]$ and $T_{k}$ the tangent plane $T_{p_{k}}M$, we assume further that the limits $\ell=\lim\ell_{k}$ and $T=\lim T_{k}$ both exist.
Let $\ell_{k}(\mathbb{C})$ be the secant line $\iota(p_{k}),\iota(q_{k})]$ in $\mathbb{C}^{n}$ and let $T_{k}(\mathbb{C})$ be the tangent space $T_{\iota(p_{k})}M_{\mathbb{C}}$. Since $\iota(p_{k})$ and $\iota(q_{k})$ are real points for all $k$, the the linear equations defining both $\ell_{k}(\mathbb{C})$ and $T_{k}(\mathbb{C})$ as varieties are exactly the same as those defining $\ell_{k}$ and $T_{k}$, respectively. Thus, the limits $\ell(\mathbb{C})$ and $T(\mathbb{C})$ both exist because the corresponding real limits exist – one may view these as limits of real sequences inside a complex Grassmannian – and they are defined by the same algebraic equations as their counterparts $\ell$ and $T$. By definition of secant lines, the image $\iota(w)$ of any $w\in\ell$ is a real point of $\ell(\mathbb{C})$. Since the pair $(M_{\mathbb{C}},N_{\mathbb{C}})$ satisfies Condition (B) by assumption, we know that $\ell(\mathbb{C})\subset T(\mathbb{C})$, whence $\iota(w)$ must be a real point of $T(\mathbb{C})$. Since $\iota(M)$ is an open subset of $M_{\mathbb{C}}$, we have $T_{\iota(p_{k})}M_{\mathbb{C}}=T_{\iota(p_{k})}\iota(M)$ for all $k$; and by the proof of Lemma 4.1, the real points of $T_{\iota(p_{k})}\iota(M)$ are identified with $T_{q_{k}}N$. Thus, $\iota(T)$ contains all the real points of $T(\mathbb{C})$, including $\iota(w)$. Since $\iota$ is injective, we have $w\in T$ as desired. ∎
@@
From an algebraic point of view our input to both the WhitStrat and the WhitneyPolar algorithm is a list of polynomials in the ring $\mathbb{R}[x_{1},\dots,x_{n}]$ [^8] Note, in a realistic computational setting the input is likely to actually be a list of polynomials in $\mathbb{Q}[x_{1},\dots,x_{n}]$, and all subsequent algebraic operations will occur in this ring. which define the real variety $X\subset\mathbb{R}^{n}$ and the corresponding complex variety $X(\mathbb{C})\subset\mathbb{C}^{n}$; in the algorithms we then perform algebraic operations starting with the polynomials in this list. We need only show that the concrete algebraic operations performed in both the WhitStrat algorithm and the WhitneyPolar algorithm leave the coefficient field of all intermediate polynomials unchanged. To see this we list the algebraic operations performed, the conclusion then follows from an analysis of the underlying algorithms in standard references such as [9].
First we list the geometric operations which occur in the WhitStrat algorithm along with their algebraic counterparts in the usual algebraic-geometric dictionary, see e.g. [9].
  1. (1)
    Intersection of varieties, which corresponds to addition of polynomial ideals.
  2. (2)
    Computation of the singular locus of a variety, from Section 2.1 (on page 5) this corresponds to computing the radical of the polynomial ideal given by taking the appropriately sized minors of the Jacobian matrix of the defining equations.
  3. (3)
    Computing the conormal variety, from Section 2.1 (on page 5) this corresponds the computation of minors of a matrix of polynomials (most of which arise from partial derivatives), and the addition and saturation of polynomial ideals.
  4. (4)
    Computing the irreducible components of a variety, which corresponds to prime decomposition of ideals.
  5. (5)
    Computing the primary decomposition of ideals.
Addition of ideals is in effect appending polynomials to a list, which leaves the coefficient field unchanged. Polynomial differentiation and the taking of minors of polynomial matrices clearly leave the coefficient field unchanged as well. Ideal saturation as well as the computation of the radical and the prime and primary decomposition can be accomplished via Gröbner basis computation, see [9] for the saturation computation and [12] for the radical, prime and primary decomposition computations. From a standard presentation of Buchberger’s algorithm to compute Gröbner basis, e.g. [9, Chapter 2, §8], we see this leaves the coefficient field unchanged. Hence all algebraic operations in items (1)–(5) leave the coefficient field unchanged and the conclusion follows.
We note that the list of algebraic options above includes computing the radical of an ideal, so for the purposes of this argument it does not matter if the input list of polynomials defines a radical ideal or not.
Similarly, the WhitneyPolar algorithm performs items (1), (2). The additional operations are:
  1. (6)
    Equidimensional decomposition of a variety (or we could use irreducible decomposition in the relevant lines, in which case this is covered by (5)).
  2. (7)
    The computation of polar varieties; by equation (8) in Section 3.2 this corresponds the computation of minors of a matrix of polynomials (most of which arise from partial derivatives), and the addition and saturation of polynomial ideals.
It can be seen from [12] that equidimensional decomposition leaves the coefficient field unchanged and, as discussed above, this is also true of the polynomial differentiation, the computation of minors of matrices of polynomials, and saturation of polynomial ideals. The conclusion follows. ∎
@@ 4.3. Dominant Morphisms between Equidimensional Varieties
It is a classical result, see e.g. [39, Chapter II, Corollary 6.7], that if $f:X\to Y$ is finite in the above sense, then it is also a proper map for any field (see Remark 2.4 for the significance of this result in our context). Over $\mathbb{K}=\mathbb{C}$, a dominant morphism is finite if and only if it is proper.
@@
then the restriction of $f$ forms a proper map $(X(\mathbb{C})-W(\mathbb{C}))\to(Y(\mathbb{C})-V(\mathbb{C}))$ — see [32, Proposition 6.1] for details. Note that the polynomials defining $V(\mathbb{C})$ and $W(\mathbb{C})$ are real; take $V$ to be the real zero set of the polynomials defining $V(\mathbb{C})$, and similarly let $W$ be the real zero set of the polynomials defining $W(\mathbb{C})$. It follows immediately that the restriction of $f$ to the difference (XW)(X-W) constitutes a proper map to the difference (YV)(Y-V).
@@
Since the stratification $Y_{\bullet}$ of $Y$ is subordinate to $V_{\bullet}$, for each stratum $R\subset Y$ we have that $R\subset V_{i}-V_{i-1}$ for some $i$, and also since $X_{\bullet}$ is subordinate to $W_{\bullet}$ we also have $f^{-1}(R)\subset W_{i}-W_{i-1}$. If $\dim(V_{i})=\dim(W_{i})$ holds, then the map $f:(W_{i}-W_{i-1})\to(V_{i}-V_{i-1})$ is proper. An appeal to the semialgebraic version of Thom’s first isotopy lemma [8, Theorem 1] achieves the desired result. ∎
@@ 5. Full Semialgebraic Stratifications
Here the $g_{i}$’s are a finite collection of polynomials in $\mathbb{R}[x_{1},\ldots,x_{n}]$. By convention, when the number of inequalities $k$ equals zero, we have $C=\mathbb{R}^{n}$. Thus, every algebraic variety is automatically a semialgebraic set in the above sense. The sets $X$ and $C$ are not uniquely determined for a given $B$ in general — we may, for instance, safely remove any polynomial generator $f:\mathbb{R}^{n}\to\mathbb{R}$ from the defining ideal of $X$ while adding $f\geq 0$ and $f\geq 0$ to the inequality locus. It is therefore customary to omit $X$ entirely and simply define $B$ as the set of points which satisfy a collection of polynomial inequalities. We find it convenient to write $B=X\cap C$ here because this allows us to isolate a relevant sub-class of semialgebraic sets.
@@
The following example highlights an important class of full semialgebraic sets which arise in the Morse theoretic study of complex analytic varieties — see [17, II.2].
@@
Whitney showed in [46] that every complex analytic variety $W\subset\mathbb{C}^{n}$ admits a Whitney stratification $W_{\bullet}$ where each intermediate $W_{i}$ is a complex analytic subvariety. Associated to each $i$-stratum $M\subset W_{i}-W_{i-1}$ is a stratified space $\mathscr{L}=\mathscr{L}_{M}$, which is called the complex link of $M$ and defined as follows. Consider a point $p\in M$ and let $A\subset\mathbb{C}^{n}$ be a generic (ni)(n-i)-dimensional affine subspace which intersects $W$ transversely at a point $p\in M$. Let $B_{\epsilon}(p)$ denote the disk containing all points within a sufficiently small radius $\epsilon>0$ around $p$ with respect to the standard euclidean distance in $\mathbb{C}^{n}$. Finally, let $\pi:A\to\mathbb{C}$ be a generic linear form sending $p$ to zero. For some positive $\delta\ll\epsilon$, one defines
@@
Since $B$ is full, we know that the interior $C^{\circ}$ of its inequality locus is a smooth open $n$-dimensional submanifold of $\mathbb{R}^{n}$. This means any open neighborhood of $X$, which is contained in $C^{\circ}$, inside $\mathbb{R}^{n}$ is identical to the same neighborhood of $X$ inside $C^{\circ}$ and therefore, the intersections $X_{i}\cap C^{\circ}$ form a Whitney stratification $X^{\prime}_{\bullet}$ of $X\cap C^{\circ}$. Alternatively the fact that $X_{i}\cap C^{\circ}$ forms a Whitney stratification $X^{\prime}_{\bullet}$ of $X\cap C^{\circ}$ follows since $C^{\circ}$ is full dimensional, meaning the intersection $X\cap C^{\circ}$ is transverse and hence the conclusion follows because the Whitney stratification of two spaces which intersect transversely is given by their intersection, [7], see also [17, Chapter 1, §1.2, (5)].
@@
Using Theorem 5.4 in practice on a full semialgebraic set $B=X\cap C$ requires constructing a stratification of the real variety $X\cap Y_{C}$; this is somewhat unsatisfactory from an efficiency perspective, since $X$ and $Y_{C}$ may intersect far away from $B$ — returning to Example 5.3, note that $P_{1}$ and $P_{2}$ might intersect $L_{2}$ at points far from $B$. We are not aware of any existing technique which can easily overcome such suboptimality.
6. Performance Analysis and Comparison
In this section we illustrate the performance of the WhitneyPolar algorithm of Section 3. The results have been collected in Table 1.
6.1. Other Methods